for what values of a and b lim x?8 ((x²+1)÷(x+1))-ax-b =0

Grade:11

3 Answers

Sher Mohammad IIT Delhi
10 years ago
limx--

= x-ax-b
if a=1 and b=0 this will be zero , and hence a=1 and b=0 is answer.

Sher Mohammad IIT Delhi
10 years ago
limx--

=x-ax-b
if a=1 and b=0 this will be zero , and hence a=1 and b=0 is answer.

Sher Mohammad IIT Delhi
10 years ago
limx--

=limx-ax-b
if a=1 and b=0 this will be zero , and hence a=1 and b=0 is answer

Sher Mohammd,
B.tech IIT Delhi